You are told that at the known positions x1 and x2 ,an oscillating mass m has speeds v1 and v2 respectively. What are the amplitude and angular frequency of the oscillations? (Hint: x(t) = B1cos(wt) + B2sin(wt))
Answer #1

X1=B1Cos(Wt1)+B2Sin(Wt1)--------------------1

X2=B1Cos(Wt2)+B2sin(Wt2)--------------------2

Differentiating With respect to t

V1=dX1/dt =-B1WSin(Wt1)+B2WCos(Wt1)----------------3

V2 =dX2/dt =-B1WSin(Wt2)+B2WCos(Wt2)----------------4

(W*1)2+32

X12W2+V12=W2*(B1Cos(Wt1)+B2Sin(Wt1))2+(-B1WSin(Wt1)+B2WCos(Wt1))2

X12W2+V12=B12W2+B22W2--------------5

(W*2)2+42

X22W2+V22=[W*(B1Cos(Wt2)+B2sin(Wt2))]2+[-B1WSin(Wt2)+B2WCos(Wt2)]2

X22W2+V22=B12W2+B22W2-----------------------6

5-6

X12W2+V12-X22W2-V22=0

(X12-X22)W2 =(V22-V12)

so angular freqeuncy

\omega =\sqrt{\frac{V_{2}^{2}-V_{1}^{2}}{x_{1}^{2}-x_{2}^{2}}}

substituting W in 5 we get

X_{1}^{2}\left ( \sqrt{\frac{V_{2}^{2}-V_{1}^{2}}{x_{1}^{2}-x_{2}^{2}}} \right )^{2}+V_{1}^{2} = (B_{1}^{2}+B_{2}^{2}) \times \left ( \sqrt{\frac{V_{2}^{2}-V_{1}^{2}}{x_{1}^{2}-x_{2}^{2}}} \right )^{2}

So amplitude

A=B_{1}^{2}+B_{2}^{2}=\sqrt{\frac{x_{1}^{2}V_{2}^{2}-x_{2}^{2}V_{1}^{2}}{V_{2}^{2}-V_{1}^{2}}}